OpenStudy (anonymous):

How do you solve an absolute value inequality? |2x|<=|x-3|

OpenStudy (anonymous):

x -3 ≥ 2x Can you solve it?

OpenStudy (anonymous):

I solved both \[2x \le x-3\] and \[2x \ge -(x-3)\]

OpenStudy (anonymous):

Exactly right :) Absolute value always yields 2 results

OpenStudy (anonymous):

However when I check my answers, \[x \le-3 \] and \[x \ge 1\] Everything get's all weird because I know that -2 works as a solution, and 0 works too.

OpenStudy (anonymous):

You check the result x = -2?

OpenStudy (anonymous):

I did, it turns out to be \[4\le5\]

OpenStudy (anonymous):

Then double check your post!

OpenStudy (anonymous):

but my requirements after I solved were that \[x \le -3\] and x being -2 doesn't fulfill the requirement.

OpenStudy (anonymous):

Your solving is correct, but it's likely that you write the question wrong somewhere :(

OpenStudy (anonymous):

It seems to me that the question is: | 2x | ≥ | x -3 | => The result is x ≥ -3 and x ≤ 1 Then your book solution is match with the interval -3 ≤ x ≤ 1

Directrix (directrix):

In interval notation, I got a solution of [-3, 1 ]. -2 is contained in that interval.

OpenStudy (anonymous):

HELLO... YOU CAN WRITE\[4x ^{2}\le(x-3)^{2}\]